OpenStudy (anonymous):

Find the equation of the line tangent to f(x) = 3x^2+ 2x + 7 at the point (1, 12).

OpenStudy (anonymous):

@Loser66

OpenStudy (abb0t):

Take the derivative of the given function. Start with that.

OpenStudy (abb0t):

The deriviative of the the given function is your slope, m. Then, plug it into the equation you know from basic algebra: \(\sf \color{red}{(y-y_1)=m(x-x_1)}\) \(\Rightarrow\) \(\sf \color{blue}{y=mx+b}\)

OpenStudy (anonymous):

ok 6x+2

OpenStudy (anonymous):

so plug in the values since we are finding it at 1,12 we do y=(1)6+2

OpenStudy (anonymous):

thats 8 so y=8x+4 is it?